OpenStudy (anonymous):

i will give medal.. Obtain the general solution of differential equation and reverse the process. xcos^2ydx + tany dy =0 x^2yy' = e^y y' = ysecx

OpenStudy (perl):

are these three seperate problems?

OpenStudy (anonymous):

yes..

OpenStudy (perl):

$$ \Large{ x\cos^2y ~dx + \tan y ~dy =0 \\~\\ \frac{ x\cos^2y ~dx + \tan y ~dy}{dx} = \frac{0}{dx} \\ x\cos^2y + \tan y ~\frac{dy}{dx} =0 \\ \tan y ~\frac{dy}{dx} =- x\cos^2y \\~\\ \frac{\tan y}{\cos^2y } ~\frac{dy}{dx} =- x \\~\\ \frac{\tan y}{\cos^2y }~ dy=- x ~dx \\~\\ \int \frac{\tan y}{\cos^2y }~ dy=\int -x ~dx } $$

OpenStudy (perl):

can you integrate that, you might want to simplify it first

OpenStudy (anonymous):

i don't know how to integrate the firsrt variable..

OpenStudy (perl):

$$ \Large{ x\cos^2y ~dx + \tan y ~dy =0 \\~\\ \frac{ x\cos^2y ~dx + \tan y ~dy}{dx} = \frac{0}{dx} \\ x\cos^2y + \tan y ~\frac{dy}{dx} =0 \\ \tan y ~\frac{dy}{dx} =- x\cos^2y \\~\\ \frac{\tan y}{\cos^2y } ~\frac{dy}{dx} =- x \\~\\ \frac{\tan y}{\cos^2y }~ dy=- x ~dx \\~\\ \int \frac{\tan y}{\cos^2y }~ dy=\int -x ~dx \\~\\ \int \frac{\frac{\sin y}{\cos y}}{ \cos^2y }~ dy=\int -x ~dx \\~\\ \int \frac{\ \sin y}{ \cos^3y }~ dy=\int -x ~dx } $$

OpenStudy (perl):

you can do u-substitution at this point

OpenStudy (anonymous):

what do you mean>>??

OpenStudy (perl):

let u = cos y , du = -sin y dy

OpenStudy (perl):

Does that look familiar?

OpenStudy (anonymous):

yes..

OpenStudy (perl):

$$ \Large{ x\cos^2y ~dx + \tan y ~dy =0 \\~\\ \frac{ x\cos^2y ~dx + \tan y ~dy}{dx} = \frac{0}{dx} \\ x\cos^2y + \tan y ~\frac{dy}{dx} =0 \\ \tan y ~\frac{dy}{dx} =- x\cos^2y \\~\\ \frac{\tan y}{\cos^2y } ~\frac{dy}{dx} =- x \\~\\ \frac{\tan y}{\cos^2y }~ dy=- x ~dx \\~\\ \int \frac{\tan y}{\cos^2y }~ dy=\int -x ~dx \\~\\ \int \frac{\frac{\sin y}{\cos y}}{ \cos^2y }~ dy=\int -x ~dx \\~\\ \int \frac{ \sin y}{ \cos^3y }~ dy=\int -x ~dx \\~\\ u = \cos y , ~ du = -\sin y~ dy, \\~\\\frac{-1}{-1} \int \frac{ \sin y}{ \cos^3y }~ dy=\int -x ~dx \\~\\\frac{1}{-1} \int \frac{- \sin y ~dy}{ \cos^3y }=\int -x ~dx \\~\\\frac{1}{-1} \int \frac{du}{ u^3 }=\int -x ~dx }$$

OpenStudy (perl):

note that -1/ -1 = 1 , so i did not change the integral . just prepared it for substitution

OpenStudy (anonymous):

okay...then??

OpenStudy (perl):

$$ \Large{ x\cos^2y ~dx + \tan y ~dy =0 \\~\\ \frac{ x\cos^2y ~dx + \tan y ~dy}{dx} = \frac{0}{dx} \\ x\cos^2y + \tan y ~\frac{dy}{dx} =0 \\ \tan y ~\frac{dy}{dx} =- x\cos^2y \\~\\ \frac{\tan y}{\cos^2y } ~\frac{dy}{dx} =- x \\~\\ \frac{\tan y}{\cos^2y }~ dy=- x ~dx \\~\\ \int \frac{\tan y}{\cos^2y }~ dy=\int -x ~dx \\~\\ \int \frac{\frac{\sin y}{\cos y}}{ \cos^2y }~ dy=\int -x ~dx \\~\\ \int \frac{ \sin y}{ \cos^3y }~ dy=\int -x ~dx \\~\\ u = \cos y , ~ du = -\sin y~ dy, \\~\\\frac{-1}{-1} \int \frac{ \sin y}{ \cos^3y }~ dy=\int -x ~dx \\~\\\frac{1}{-1} \int \frac{- \sin y ~dy}{ \cos^3y }=\int -x ~dx \\~\\\frac{1}{-1} \int \frac{du}{ u^3 }=\int -x ~dx \\~\\\frac{1}{-1} \int u^{-3}~du=\int -x ~dx \\~\\\frac{1}{-1}\cdot \frac{u^{-3+1}}{(-3+1)}= -\frac{x^2}{2}+C \\~\\-1\cdot \frac{u^{-2}}{(-2)}= -\frac{x^2}{2}+C \\~\\ \frac{(\cos y)^{-2}}{2}= -\frac{x^2}{2}+C \\~\\ \frac{1}{2 \cos^2 y }= -\frac{x^2}{2}+C }$$

OpenStudy (perl):

That is the general solution. To reverse the process i think they mean, take the derivative

OpenStudy (perl):

thats a good way to check that you integrated correctly, by the way

OpenStudy (perl):

$$\Large{ \\~\\ \frac{(\cos y)^{-2}}{2}= -\frac{x^2}{2}+C \\~\\ 2\cdot \frac{(\cos y)^{-2}}{2}= 2\cdot \left(-\frac{x^2}{2}+C \right) \\~\\ (\cos y)^{-2}= -x^2+2C \\~\\ \rm take~ implicit~ derivative ~ wrt ~ x \\~\\ \frac{d}{dx} (\cos y)^{-2}= \frac{d}{dx} \left(-x^2+2C \right) \\~\\ -2 (\cos y)^{-3}\cdot \frac{dy}{dx} = -2x + 0 } $$

OpenStudy (perl):

now that doesnt look like the original differential equation, so we have to look at our original differential equation. can we bridge these two expressions

OpenStudy (perl):

by algebra and trig manipulation

OpenStudy (perl):

$$ \Large{ \\~\\ \frac{(\cos y)^{-2}}{2}= -\frac{x^2}{2}+C \\~\\ 2\cdot \frac{(\cos y)^{-2}}{2}= 2\cdot \left(-\frac{x^2}{2}+C \right) \\~\\ (\cos y)^{-2}= -x^2+2C \\~\\ \rm take~ implicit~ derivative ~ wrt ~ x \\~\\ \frac{d}{dx} (\cos y)^{-2}= \frac{d}{dx} \left(-x^2+2C \right) \\~\\ -2 (\cos y)^{-3}\cdot \frac{dy}{dx} = -2x + 0 \\~\\ (\cos y)^{-3}\cdot \frac{dy}{dx} = x \\~\\ \frac{1}{(\cos y)^{3}}\cdot \frac{dy}{dx} = x \\~\\ \frac{1}{\cos^2 y }\cdot \frac{1}{\cos y}\cdot \frac{dy}{dx} = x \\~\\ \frac{1}{\cos y}\cdot \frac{dy}{dx} = x\cdot \cos^2 y \\~\\ \frac{\sin y}{\sin y}\cdot \frac{1}{\cos y}\cdot \frac{dy}{dx} = x\cdot \cos^2 y }$$

OpenStudy (anonymous):

siny/siny?? whre it came from??

OpenStudy (perl):

i introduced it, but it might not help

OpenStudy (anonymous):

1-sin^2y = cos^2y??

OpenStudy (perl):

oh i made a mistake when i took derivative

OpenStudy (perl):

grrrr

OpenStudy (perl):

$$ \Large{ \\~\\ \frac{(\cos y)^{-2}}{2}= -\frac{x^2}{2}+C \\~\\ 2\cdot \frac{(\cos y)^{-2}}{2}= 2\cdot \left(-\frac{x^2}{2}+C \right) \\~\\ (\cos y)^{-2}= -x^2+2C \\~\\ \rm take~ implicit~ derivative ~ wrt ~ x \\~\\ \frac{d}{dx} (\cos y)^{-2}= \frac{d}{dx} \left(-x^2+2C \right) \\~\\ -2 (\cos y)^{-3}\cdot \color{red}{ (-\sin y )} \frac{dy}{dx} = -2x + 0 } $$

OpenStudy (anonymous):

it's okY

OpenStudy (anonymous):

COS^2y * COSy ? so that sin y/cos y will become tany...

OpenStudy (anonymous):

am i right?

OpenStudy (perl):

$$ \Large{ \\~\\ \frac{(\cos y)^{-2}}{2}= -\frac{x^2}{2}+C \\~\\ 2\cdot \frac{(\cos y)^{-2}}{2}= 2\cdot \left(-\frac{x^2}{2}+C \right) \\~\\ (\cos y)^{-2}= -x^2+2C \\~\\ \rm take~ implicit~ derivative ~ wrt ~ x \\~\\ \frac{d}{dx} (\cos y)^{-2}= \frac{d}{dx} \left(-x^2+2C \right) \\~\\ -2 (\cos y)^{-3}\cdot \color{red}{ (-\sin y )} \frac{dy}{dx} = -2x + 0 \\~\\ 2 (\cos y)^{-3}\cdot \color{red}{ (\sin y )} \frac{dy}{dx} = -2x \\~\\ (\cos y)^{-3}\cdot \color{red}{ (\sin y )} \frac{dy}{dx} = -x \\~\\ \frac{1}{(\cos y)^{3}}\cdot \color{red}{ (\sin y )} \frac{dy}{dx} = -x \\~\\ \frac{1}{(\cos y)^{2}}\cdot \frac{1}{\cos y } (\sin y ) \frac{dy}{dx} = -x \\~\\ \frac{1}{(\cos y)^{2}}\cdot \frac{\sin y}{\cos y } \cdot \frac{dy}{dx} = -x } $$

OpenStudy (perl):

right, see if you can finish it ☺